Fermi National Laboratory


Accelerator Update

Monday April 12
The day shift began with Operations monitoring the Recycler stash, with the Tevatron (TeV) recovering from its quench, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiment taking beam.

At 9:40 AM, Operations put the TeV through a wet squeeze and then began a study period.

Operations turned off the MCenter beam at 10:44 AM for Main Injector (MI) slow accelerations studies.

The TeV study beam aborted at sector A1 due to a tuning problem at 11:44 AM.

A Pbar expert began a ten-minute lithium lens target study at noon.

MI and Pbar began its first transfer of antiprotons for studies at 1:31 PM. The last of ten transfers occurred at 3:20 PM. Unfortunately, MI couldn't accelerate the beam to the 2.5Mhz needed for the study.

At 2:15 PM, Operations put the TeV through several wet squeezes for orbit studies.

Operations reestablished beam to the MCenter experiment at 5:51 PM.

At 7:19 PM, A Booster kicker (MKS03) tripped off and wouldn't reset. This shut off Switchyard (SY) beam. Operations called in an expert and he had the kicker repaired by 8:56 PM.

A Pbar expert began a forty-five minute Debuncher momentum cooling study at 9:12 PM. Pbar then resumed stacking.

A Booster RF station (BRF16) tripped off at 9:35 PM and wouldn't reset. Operators bypassed this station for about 30 minutes and then turned it back on with no problem.

Tuesday April 13
The midnight shift began with Pbar stacking, with TeV experts conducting studies, and with MiniBooNE and MCenter taking beam.

At 6:58 AM, two Booster RF stations (BRF13 & 14) tripped off due to a vacuum burst. Both stations reset okay.

TeV experts paused their studies at 9:10 AM for an hour to allow the Collider Detector at Fermilab (CDF) and D-zero (D0) experimenters access to their collision halls.

Technicians needed to work on some SY correction magnets located in TeV sector F. Operators made the area ready for an access at 9:22 AM. This access shuts off beam to MiniBooNE and SY.

Both CDF and D0 completed their accesses by 10:53 AM. The TeV access ended at 11 AM.

Operations reestablished stacking at 11:30 AM and had beam to MiniBooNE by 11:35 AM.

Operations reestablished beam to MCenter at noon.

Operations established a small proton only study-store at 1:21 PM.

A TeV RF station (TRF7) tripped off at 2:22 PM due to what appeared to be a water heater ground fault. Operators found out that a tech was taking some readings and inadvertently tripped the station.

At 3:22 PM, a TeV quadrupole magnet tripped off its bulk power supply on a quench indication.

Pbar resumed stacking at 3:24 PM.

Operations turned off Meson beam at 3:40 PM so that MCenter could access their enclosures.

Operations put the TeV into shot setup at 4:31 PM.

At 7:35 PM, Operations established store 3399 with an initial luminosity of 42.81E30. Pbar resumed stacking.

A TeV expert successfully performed a collimator move at 10:16 PM.

Wednesday April 14
The midnight shift began with Operations monitoring store 3399 and a Recycler stash, with Pbar stacking, and with MiniBooNE taking beam.

At 7:57 AM, Operators noted a problem with the TeV tune. Operators investigated, but found nothing wrong. Their investigation continued.

Plans for Today
The plans for today are to allow Pbar to conduct studies this afternoon and then to establish a new TeV store tonight. Tomorrow (Thursday, April 15th), Recycler will conduct a twelve-hour access to check for vacuum leaks.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/14/2004   email Fermilab